OpenStudy (mpj4):

How to find the area of the circle that's not inside the triangle with radius of 8 units and a cylic triangle of angles 60, 70 & 50.

OpenStudy (anonymous):

is this a quiz test?

OpenStudy (mpj4):



OpenStudy (mpj4):

just a book question.

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

do you know the answer? please teach me i dont know

OpenStudy (mpj4):

It's 121.26 units squared but it came from the book.

OpenStudy (mpj4):

anyone? D:

OpenStudy (dumbcow):

 You can find sides of triangle using Law of cosines \[a^2 = 8^2 + 8^2 - 2(8)(8) \cos \theta\] you can find heights of each smaller triangle using sine \[h = 8 \sin \theta\]

OpenStudy (mpj4):

Thank you sir, I did that a few minutes ago since it took too long for someone to respond xD, i found that the area of the triangle was 79.80 and the area of the circle was 201.06, subtracting the former with the latter gives me 121.26 units squared.

OpenStudy (mpj4):

lol, messed up with former and latter, i meant latter with former xD

OpenStudy (dumbcow):

:)